June 6, 2013 Solenoid Meeting
Thursday June 6,2013 at 10am At Jefferson Lab, the meeting will be held in room ARC 231
E. Wolin, Yi Quiang, J. Creel, E. Chudakov, C. Rode, T. Whitlatch
- Discuss last weeks minutes https://halldweb.jlab.org/wiki/index.php/May_28,_2013_Solenoid_Meeting
- Work accomplished since last meeting
- 15A test minus quench detector checks (magnet still resistive)
- Smooth ramp appeared to work ok
- Balanced GFD for 16ohm short
- Preliminary 140A test procedure written Media:140A procedure rev C.pdf
- Discuss path forward plan Media:wSchneider path forward.pdf
- Probirs margin comparison Media:Margin comparison.pdf
- Refrigerator status
- Do we go right into high current from 140A if ready?
- Path Forward .......
- Documents available at M:\halld_engineering\Solenoid\Commissioning\Response to Quench
We went through the agenda discussing last week's meeting first.
The 15A test plan was implemented through this past Monday and most items were completed with satisfactory results. Portions of the 15A test plan was not completed due to the fact that the magnet was still dominated by resistive load since the magnet is not yet at superconducting temperatures. The magnet was only brought up to 2+ amps and then brought back down. Voltage spikes were still seen on ramp up but the current ramp appeared smooth. The hardware quench detector still needed to be over-ridden when the power supply was turned on and ramped to 2+ amps and then fast dumped. This was the maximum current applied with quench detector over-ridden. It was decided that no more useful testing could be done until the magnet is at a stable, superconducting temperature. Therefore, the balancing of the quench detector and the checking for turn to turn shorts portion of the 15A test will be moved to the next set of testing when superconducting. We also looked through Bill Schneider's "Path forward" and are in general agreement with this and the items discussed below should address the items outlined.
- The initial 140A test procedure has been written and distributed by George and awaiting additional comments. This part basically checks all the systems required during high current runs as well as the turn to turn short check and QD balancing.
- We discussed Probir's comparison of margins using different conditions. All margins are based on the spec Alcorn used and not the short samples. From this, it looks like subcoil 1A had a margin of ~0.5K 1500A/4.6K. Using 1300A at 4.6K or 1350A at 4.5K will give a margin of ~1K.
- Discussed both the hardware and PLC QDs. The known inductances from the previous 1500A run up will be used to program the PLC QD to stay balanced throughout the current ramp up. A "dial potentiometer" will be installed onto channel 3 on the hardware QD to be able to manually balance (with precision) as the current is raised. - Elliot will follow both items through.
- Danfysik has finally returned a quote for engineering and parts for the balancing of the power supply for 25H. The cost is $3.6K and the PR is already in the system. Giles will follow through on getting this implemented.
- George will add a check of the resistance of the full magnet (minus the leads when superconducting)and acceptable readings to the procedure.
- After a discussion of the "buggers" (items that are changed from normal operating conditions), it was agreed that all known buggers should be put into the procedure and approved up front. Anything added later would have to be approved by the responsible engineer and/or the DSO.
- The refrigerator is still being worked on and Jonathan thinks it will be ready by the end of next week. Carbon and oil are being cleaned out of the system. The question remains on when to start cooldown of the magnet. Consensus is that we should wait until at least the 140A portion of the procedure is signed off and there is general consensus that we will go right into the high current testing if the 140A testing is successful. By successful, it is meant that all critical systems are working and the quench detector operates without false trips. Jonathan is concerned that this may be our last opportunity to get the magnet cold without a major overhaul to the refrigerator.
1. Incorporate quench detection balance/tests into the 140A procedure (Rev C) - George
2. Update hardware QD with dial pot and PLC with inductance algorithm - Elliot
3. Add full resistance measurement and acceptable value to 140A procedure - George
4. Identify all known buggers in the procedure for approval - George
5. Clean up refrigerator and get ready for cooldown - Jonathan
6. Determine if turn to turn short could have caused quench in coil 1A - Eugene
7. Determine what is a safe current to run at using the existing refrigerator - all
8. Set up a meeting with all parties to get agreement on the path forward - Tim
9. Write up determination of what it would take to catch a quench once it starts - Probir/George
10. Write up final analysis of quench (suppositions) - George/Eugene
11. Write up high current portion of procedure (Rev D) - George
12. Balance Power supply for 25H - Giles
13. Order 2 additional burst discs - Tim
1. Complete the 140A procedure including the quench balance and turn to turn short check not completed in the 15A test (to be completed and signed off by June 14)
2. Continue repair of refrigerator and fill magnet with liquid (to be completed by June 19)
3. Complete and sign off maximum current plan (to be completed by June 19)
4. Perform 140A testing ( to be completed by June 26)
5. Perform High current commissioning ( to be completed by June 28)
6. Map magnet (to be completed by July 3)